# CasTex: Cascaded Text-to-Texture Synthesis via Explicit Texture Maps and Physically-Based Shading

### v1

@@ 4.1 Limitations of Score Distillation Sampling for Latent Diffusion Models
Following the common practice, we convey the analysis on a task of image generation. Specifically, we use a mapping $g_{id}(\theta,c)=\theta$ that outputs a 2D image $\theta$ and does not rely on the second stochastic variable. The goal is to generate images on par with the diffusion model samples with a gradient descent algorithm. Intuitively, the SDS gradients guide images $g(\theta,c)$ towards the manifold of natural images. Using the chain rule, these gradients also propagate to the parameter $\theta$. However, in the case of latent diffusion models guidance happens within the latent space: $g(\theta,c)=\text{enc}(g_{id}(\theta,c))$. As a result, the training signal must propagate through the encoder.
Next, to mitigate some of the potential issues of SDS in our analysis, we consider a toy diffusion model that represents a degenerate distribution $\delta(x-x^{*})$ concentrated at $x^{*}$. The problems may include a poorly trained score estimation model or the excessive variance of the stochastic gradient 1. For the degenerate distribution, we use the optimal model that separates the Gaussian noise from the original sample
@@
Note that for the model from Eq. 3 and constant weights $w(t)=1$ the gradient in Eq. 1 simplifies to the deterministic gradient of the mean squared error:
@@
In other words, in the specified toy setup, we can assume
@@
To analyze SDS, we ran the optimization in various setups and report the results in Figure 2. For the pixel-space diffusion model with $g(\theta)=\theta$, gradient descent rapidly converges to $\theta\approx x^{*}$ (top row, left). For the latent-space model we assume $x^{*}$ to be the latent code of the same image and minimize $\|\text{enc}(g(\theta))-x^{*}\|^{2}$. Although gradient descent converges and the decoder maps $\text{enc}(\theta)$ to the original image (top row, right), the optimized parameter $\theta$ itself is a poor approximation to the original image (top row, middle). In practice, SDS with latent diffusion models finds globally optimal solutions that fall outside of the natural image domain.
Notably, the above observation does not limit text-to-3D applications. However, in the above toy setup and texture synthesis domain the observed artifacts limit the applicability of SDS. To avoid the artifacts, a recent work [41] proposed to parameterize textures with a deep convolutional network. Their solution builds upon the observations of [38] and acts as an implicit regularizer. It helps avoiding solutions outside of the natural image domain, but requires running an additional convolutional and, as we show below, still leads to noticeable image artifacts.
In the bottom row of Figure 2, we report SDS performance with an implicit image parameterization from [41]. We denote the parameterization as DIP (Deep Image Prior), following [38]. In this case, the image is an output of a UNet network with weights $\theta_{dip}$ and a fixed input sampled from the standard Gaussian distribution. In the context of the SDS, we can interpret the forward pass of the UNet as the differentiable rendering algorithm $g_{\text{UNet}}(\theta_{dip})$ constrained to generate natural images.
We minimize Eq. 5 for the implicit parameterization assuming both latent and pixel-space models. The implicit parameterization significantly improves the result for the latent diffusion model (middle row). The artifacts present in the explicit parameterization are less noticeable. Still, even with the implicit parameterization SDS in the latent space is unable to match the reconstruction quality of the reconstructed latent code (middle and right bottom row images). In our preliminary experiments, we tried to impose additional regularization on $\theta$ using the decoder outputs, but did not achieve noticeable improvements compared to the default SDS. Besides that, pixel-space optimization achieves significantly better result for both explicit and implicit image parameterization.
Given the above observations, in our work, we use SDS with pixel-space models.
4.2 Applying SDS with Super-Resolution Diffusion Models
The downside of pixel-space diffusion models is that the model for high-resolution images require significant resources. While latent diffusion models employ a VAE [18] to encode a high-resolution image into a low-dimensional latent code, the typical solution for pixel-space models is to generate an image in a cascaded fashion. At the first stage, diffusion model generates a low-resolution image. During the second stage, a super-resolution model generates a high-resolution version of the image from the first stage.
In [29], the authors run SDS only with a low-resolution model. However, the low-resolution model may be unable to recover fine-grained details and high-frequency features as its training domain was limited to down-scaled images. To achieve highly detailed textures, in our work we mimic the architecture of cascaded diffusion models. First, we run SDS with a low-resolution diffusion model. Then, we refine the output of the first stage with SDS combined with a super-resolution model.
Figure 3: Score distillation sampling applied to a super-resolution diffusion model. Horizontal axis represents progress over time. Top row: the same image $g(\theta)$ used both as a diffusion model input $x$ and the condition $x_{cond}$ for super-resolution. The resulting procedure is unstable and significantly diverges from the original image. Bottom row: same procedure with fixed super-resolution condition $x_{cond}=g(\theta_{0})$. The condition anchors the dynamics to the initial image.
In text-to-image setup, the second stage model takes a low-resolution image as an input along with a text prompt and time. This leads to the two natural ways of applying SDS to refine the first stage output. The first option is to pass the rendered image $g(\theta,c)$ as the diffusion model argument in SDS and a down-scaled copy of the image as a condition. Intuitively, in this case, the algorithm gradually updates the texture by adding the details lacking after the previous gradient step. The second option is to freeze the parameter $\theta_{0}=\theta$ obtained at the first stage and run SDS using a copy of $\theta$ as the new parameter and the previous $\theta_{0}$ as a condition. In this case, the first stage output in the condition acts as an anchor for the gradient descent dynamics.
In our experience, the procedure in the first solution tends to be less stable and leads toward out-of-domain images, whereas the second approach tends to improve the resolution while preserving the details of the first stage result. We illustrate our observations in Figure 3.
In summary, we propose a texture synthesis approach based on a cascaded modification of the SDS. Figure 1 is a schematic representation of the overall approach.
@@ 5.1 Evaluation setup
We compared the proposed method with three baselines that represent various approaches to texture synthesis. First, Text2Tex [6] is a well-established back-projection-based method generating non-PBR textures. Second, Paint-it [41] is a simplistic optimization-based method using SDS. The crucial distinction between our method and Paint-it is the underlying diffusion model and texture parameterization. Finally, FlashTex [13] is another recent approach using SDS. It addresses the limitations of previous methods by fine-tuning the model used for guidance on multi-view 3D data with controlled lighting. Again, all of the baselines are based on latent diffusion models for texture synthesis.
@@ 5.2 Main Results
Figure 4: Qualitative comparison. We synthesized textures for models from the Objaverse dataset using the proposed approach and a number of recent competing methods. Our method generates seamless textures with softer colors compared with latent diffusion-based approaches.
@@
Size FID KID Runtime
($\downarrow$) ($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
Text2Tex 2.5B 21.3 4.1 549
@@
CasTex (M) 400M 24.9 5.6 45
CasTex (M+L) 1.6B 23.2 5.0 628
CasTex (XL) 4.3B 22.8 4.4 174
CasTex (XL+L) 5.5B 20.5 3.1 710
@@
Automated quality assessment. We report FID and KID scores in Table 1. Our smallest single-stage setup outperformed the two optimization-based approaches. The two-stage cascaded diffusion approach achieved the best performance, outperforming all baselines. We illustrate the two stages in Figure 6 and provide additional illustrations in Appendix C. The second stage significantly refines textures, enhancing fine-grained details and coherence.
Notably, the weak performance of FlashTex on the benchmark is the result of the limited generalization abilities of the Light ControlNet proposed by the authors. To our surprise, Text2Tex delivered competitive results. Even though the method is the oldest baseline, the back-projected photo-realistic images lead to photo-realistic textures and renders.
@@
User preference study. In addition, we conduct a human study to compare the perceptual quality of our generated textures against baseline methods. The evaluation is conducted by professional assessors, asked to make a decision between two rendered textures given a textual prompt. The decision is made according to the texture quality and textual alignment. For each pair, three responses are collected, and the final result is determined by majority voting. For evaluation, we use $256$ random objects from the previously mentioned Objaverse subset.
As shown in Figure 5, our two-stage approach outperforms all the baselines. Figure 4 presents a qualitative comparison of our method with the three baseline approaches. The comparison against our single-stage approach confirms that the second stage is highly effective for further texture refinement.
Also, we provide the results of our method using the smaller version of DeepFloyd-IF that contains 400M parameters. We observe that our method still outperforms competing works.
@@ 5.3 Ablation study
We conducted a series of experiments to understand the role of the key design choices in our approach.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
CasTex (diffuse map only) 21.8 4.0
CasTex 20.5 3.1
@@
Figure 6: Textures after the first and the second stage.
Do the PBR components improve the texture synthesis? PBR components enable the modeling of complex visual effects like reflections and material properties, avoiding baked-in lighting and improving realism.
We compare our full PBR-based approach with a baseline that generates only the diffuse map, one of the PBR components. As shown in Table 2, generating the full set of PBR maps leads to better FID and KID scores compared to generating the diffuse map alone. We provide visualization of PBR maps in Appendix E. Despite better metrics, a closer study of illustrations indicates that the method occasionally stuggles with baked-in lighting.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
Text2Tex 21.3 4.1
CasTex (Text2Tex init, $k_{d}$ frozen) 21.3 4.2
CasTex (Text2Tex init, $k_{d}$ trained) 20.5 3.9
@@
As shown in Table 3, our second stage can effectively generate PBR textures when initialized with diffuse maps from Text2Tex. We explore two configurations: (1) where we keep the initial diffuse map fixed ($k_{d}$ frozen) and (2) where we fine-tune the diffuse map during PBR generation ($k_{d}$ trained).
@@
The results show that when the diffuse map is kept frozen, the FID score remains unchanged with a slight increase in KID. This may be because freezing $k_{d}$ also preserves any baked lighting information present in the initial diffuse map, preventing our method from correcting for these lighting inconsistencies. In contrast, when we allow diffuse map optimization ($k_{d}$ trained), our approach achieves significant improvements in both metrics, yielding the best FID and KID. This demonstrates that enabling diffuse map refinement allows our method to correct lighting artifacts and better harmonize the entire PBR material set, highlighting the potential of our approach as a versatile enhancement module for various texture generation pipelines.
FID KID
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$)
Paint-it (Paint-it envlight) 25.2 6.7
Paint-it (proposed envlight) 24.5 6.4
CasTex (Paint-it envlight) 23.3 5.5
CasTex (proposed envlight) 20.5 3.1
Table 4: Ablation of environment lighting. We discovered that environmental lighting may play a crucial role in the plausibility of resulting physically-based textures. The default environmental lighting map in Paint-it presents an even gray lighting across all directions. The difference between a specular and a rough diffuse surface with such lighting is barely distinguishable. Instead, we propose an environmental lighting map with a single pronounced light source and augment lighting with horizontal rotations.
@@
How does environment lighting affect results? Among recent works, Paint-it is the most similar to ours, as it also relies solely on SDS-based optimization. Unlike our approach, their rendering algorithm uses a brighter, uniform lighting environment map, whereas our setup combines a pronounced light source and a weaker ambient lighting component.
To understand the effect of environment lighting on SDS-based optimization, we evaluate both Paint-it’s lighting setup and our alternative configuration. As shown in Table 5, Paint-it achieves better results when using our lighting setup, slightly improving its FID and KID. Similarly, our methods performance degrades with Paint-it’s environment lighting. We hypothesize that, unlike uniform lighting, a single pronounce light source allows to highlight specular components of the PBR texture during training. The hypothesis and the above observations agree with the results from [13].
FID KID Runtime
($\downarrow$) (×$\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
CasTex (DIP, XL) 23.2 4.6 229
CasTex (XL) 22.8 4.4 174
CasTex (DIP, XL+L) 21 3.8 862
CasTex (XL+L) 20.5 3.1 710
Table 5: Deep Image Prior ablation. Even though DIP plays a crucial role in texture synthesis with latent diffusion models, for the cascaded diffusion models the induced regularization is obsolete and slightly reduces the resulting scores. Also DIP-based variants are more time-consuming, while all experiments generate textures at 1024$\times$1024 resolution.
DIP versus explicit parameterization. Following the approach of our closest competitor, Paint-it, we integrate Deep Image Prior (DIP) [38] into our method and evaluate its impact on texture optimization (see Section 4.1 for details). As shown in Table 5, our DIP-based variant achieves results comparable to our standard explicit parameterization approach but performs slightly worse in both FID and KID metrics while being significantly slower. Since explicit parameterization already delivers strong results and works faster, the added complexity of DIP does not offer a clear advantage in our framework.

### v2

@@ 4 Score Distillation Sampling Analysis
Following common practice, we conduct the analysis in the simplified setting of image generation. We use a mapping $g_{id}(\theta,c)=\theta$ that produces a 2D image $\theta$ and while ignoring the stochastic variable. The goal is to recover images that are consistent with the diffusion model’s distribution using gradient descent. Intuitively, SDS gradients are expected to push generated images $g(\theta,c)$ toward the maniforld of natural images. Using the chain rule, these gradients also propagate to the parameter $\theta$. However, with latent diffusion models, guidance occurs in latent space: $g(\theta,c)=\text{enc}(g_{id}(\theta,c))$. As a result, the training signal must propagate through the encoder, placing additional constraints on the optimization trajectory.
To isolate the behavior of SDS and exclude confounding factors such as model mis-specification or stochastic gradient noise, we consider a toy diffusion model that represents a degenerate distribution $\delta(x-x^{*})$ concentrated at a single image $x^{*}$. This setup removes issues arising from imperfect score estimation or excessive variance in Eq. 1. For the degenerate distribution, the optimal score model reduces to a closed-form denoiser:
@@
With constant weights $w(t)=1$, the SDS gradient simplifies to the deterministic gradient of the mean squared error:
@@
Thus, in this idealized setting we can write
@@
We optimize Eq. 5 under various parameterizations and report the results in Figure 3. For pixel-space diffusion, where $g(\theta)=\theta$, gradient descent quickly converges to $\theta\approx x^{*}$ (top row, left). For latent-space model, we assume $x^{*}$ to be the latent code of the same image and minimize $\|\text{enc}(g(\theta))-x^{*}\|^{2}$. Although the gradient descent converges and the decoder maps $\text{enc}(\theta)$ to the original image (top row, right), the optimized parameter $\theta$ itself is a poor approximation to the original image (top row, middle). In other words, SDS with latent diffusion models frequently finds near-optimal latent solutions whose pre-images fall outside the natural image domain.
This phenomenon does not significantly affect some text-to-3D tasks, where the optimization variable is not directly visualized. However, in the context of texture synthesis—where $\theta$ is the texture—these artifacts severely limit the applicability of SDS.
To address this issue, [36] proposes parameterizing textures via a deep convolutional network, following the Deep Image Prior (DIP) framework [35]. This parameterization acts as an implicit regularizer that discourages unnatural images. Although effective to some extent, it introduces an additional convolutional network into the pipeline and, as we show below, still leaves noticeable artifacts.
The bottom row of Figure 3 illustrates SDS optimization with the DIP parameterization. Let the generated image be produced by a U-Net with parameters $\theta_{dip}$ and a fixed Gaussian noise input. We denote this differentiable generator as $g_{\text{UNet}}(\theta_{dip})$. We optimize Eq. 5 under both pixel-space and latent-space models. The DIP parameterization substantially improves the solution for latent diffusion (middle row), reducing the artifacts seen with the explicit parameterization. However, even with DIP, SDS in latent space fails to match the reconstruction quality of directly optimizing the latent code (middle and right bottom row).
We also experimented with additional regularization based on decoder outputs but observed no meaningful improvement over standard SDS. In contrast, pixel-space optimization achieves significantly better reconstruction in both explicit and implicit parameterizations.
Figure 3: Optimization based image generation for a toy single-mode diffusion model. We study the effect of image parameterization and the difference between pixel and latent-space diffusion models. Score Distillation Sampling (SDS) finds the distribution mode for both latent-and pixel-space diffusion models. Top row: the pixel space model achieves near-perfect reconstruction quality measured by PSNR. The latent code $\text{enc}(\theta_{\text{RGB}})$ is also close to the mode and leads to high quality reconstruction. However, for the latent model the optimized parameter $\theta_{\text{RGB}}$ fails to approximate the original image. Bottom row: The implicit regularization proposed in [36] significantly improves the results for SDS for the latent diffusion model, but still falls short of the image reconstruction given by the decoder or pixel space diffusion model.
@@ 5.1 Evaluation setup
We compared the proposed method with four baselines that represent various approaches to texture synthesis. First, Text2Tex [6] is a well-established backprojection-based method generating non-PBR textures. Second, SyncMVD [20] is a recent method that enforces consistency by generating and synchronizing multiple views. Third, Paint-it [36] is a simplistic optimization-based method using SDS. The crucial distinction between our method and Paint-it is the underlying diffusion model and texture parameterization. Finally, FlashTex [12] is another recent approach using SDS. It addresses the limitations of previous methods by fine-tuning the model used for guidance on multiview 3D data with controlled lighting. Again, all of the baselines are based on latent diffusion models for texture synthesis.
@@ 5.2 Main Results
Figure 4: Qualitative comparison. We synthesized textures for models from the Objaverse dataset using the proposed approach and several recent competing methods. Our method generates seamless textures with softer colors compared with latent diffusion-based approaches.
@@
Method Size FID KID Runtime
($\downarrow$) ($\downarrow$) ($\times 10^{-3}$, $\downarrow$) (seconds, $\downarrow$)
Non-SDS methods
Text2Tex 2.5B 21.3 3.9 549
SyncMVD 1.5B 20.3 2.7 55
SDS methods
@@
CasTex (M) 400M 24.6 5.4 38
CasTex (M+L) 1.6B 22.9 4.6 628
CasTex (XL) 4.3B 22.4 4.2 174
CasTex (XL+L) 5.5B 19.5 3.0 710
@@
Automated quality assessment. We report FID and KID scores in Table 1. Our smallest single-stage setup outperformed the two optimization-based approaches. The two-stage cascaded diffusion approach achieved the best performance, outperforming all baselines. We illustrate the two stages contrasting with the noisy results from latent diffusion and provide additional illustrations in Appendix C. The second stage significantly refines the textures, enhancing fine-grained details and coherence.
Notably, the weak performance of FlashTex on the benchmark is the result of the limited generalization abilities of the Light ControlNet proposed by the authors. To our surprise, Text2Tex delivered competitive results. Even though the method is the oldest baseline, the back-projected photo-realistic images lead to photo-realistic renders. SyncMVD also achieved strong scores, outperforming other baselines, which highlights the effectiveness of its multi-view synchronization approach.
@@
User preference study. In addition, we conduct a human study to compare the perceptual quality of our generated textures with the baseline methods. The evaluation is conducted by professional assessors, asked to make a decision between two rendered textures given a textual prompt. The decision is made according to the texture quality and textual alignment. For each pair, three responses are collected and the final result is determined by majority vote. For evaluation, we use $256$ random objects from the previously mentioned Objaverse subset.
As shown in Figure 5, our large two-stage approach outperforms all the baselines. Figure 4 presents a qualitative comparison of our method with the four baselines. The comparison against the first stage confirms that the second stage is highly effective for further texture refinement.
In addition, we provide the results of our method using the smaller version of DeepFloyd-IF that contains 400M parameters. We observe that our method still outperforms most competing works.
@@ 5.3 Ablation study
We conducted a series of experiments to understand the role of key design choices in our approach.
Method FID KID
($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
CasTex (diffuse map only) 21.2 3.7
CasTex 19.5 2.4
@@
Figure 6: Qualitative analysis of PBR components generated by CasTex. Lighter colors indicate higher metallic and roughness values for the corresponding texutre components. The microscope shows a clear separation between painted non‑metals and polished steel optics, with low roughness for the lenses and higher values for rubber and plastics. The banjo treats the wooden shell and drumhead as non‑metal while assigning metalness to the hardware and a semi‑gloss roughness to the lacquered wood. The binoculars further demonstrate semantic consistency by distinguishing matte rubber eyecups and strap from the painted housing and metallic fasteners. Across all three, the normal map contributes fine surface detail without altering the geometry.
Do the PBR components improve textures? PBR components enable the modeling of complex visual effects like reflections and material properties, avoiding baked-in lighting and improving realism, as qualitatively demonstrated in Figure 6 (additional illustrations in Appendix E).
We compare our full PBR-based approach with a baseline that generates only the diffuse map, one of the PBR components. As shown in Table 2, generating the full set of PBR maps leads to better FID and KID scores compared to generating the diffuse map alone. Despite better metrics, a closer study of illustrations indicates that the method occasionally stuggles with baked-in lighting.
Method FID KID
($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
Text2Tex 21.3 3.9
CasTex (Text2Tex init, $k_{d}$ frozen) 20.9 3.7
CasTex (Text2Tex init, $k_{d}$ trained) 19.7 3.6
@@
As shown in Table 3, our second stage can effectively generate PBR textures when initialized with diffuse maps from Text2Tex. We explore two configurations: (1) keeping the initial diffuse map fixed ($k_{d}$ frozen) and (2) fine-tuning the diffuse map during PBR generation ($k_{d}$ trained).
@@
The results show that even when the diffuse map is kept frozen, our approach improves both FID and KID scores over the original Text2Tex output. This improvement is likely limited because freezing $k_{d}$ also preserves any baked-in lighting information present in the initial diffuse map, preventing our method from correcting these inconsistencies. In contrast, when we allow diffuse map optimization ($k_{d}$ trained), our approach achieves a significant improvement in FID and a further, albeit smaller, improvement in KID. This shows that enabling diffuse map refinement helps to correct lighting artifacts and better harmonize the entire PBR material set, highlighting its potential as a versatile enhancement module for various texture generation pipelines.
Studio light eval Paint-It light eval
Method (Training envlight) FID KID FID KID
($\downarrow$) ($\times 10^{-3}$, $\downarrow$) ($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
Paint-It (Paint-It) 25.2 6.7 26.8 6.2
CasTex (Paint-It) 22.2 4.6 24.7 4.5
Paint-It (Proposed) 24.5 6.4 26.9 6.6
CasTex (Proposed) 19.5 2.4 24.1 4.4
Table 4: Ablation of environment lighting. We investigate how different training and evaluation environment maps affect PBR texture quality. Both CasTex and Paint-It are trained on either the Paint-It or our Proposed maps. Performance is then evaluated under two distinct lighting conditions: our proposed Studio (which differs from both training maps) and the Paint-It maps. Our full pipeline (CasTex trained with the Proposed envlight) achieves the best results in both evaluation scenarios. Notably, CasTex also outperforms Paint-It even when both are trained and evaluated on the less expressive Paint-It map.
@@
How does environment lighting affect results? Among recent works, Paint-it is the most similar to ours, as it also relies solely on SDS-based optimization. Unlike our approach, their rendering algorithm uses a brighter, uniform lighting environment map for training, which can make it difficult to distinguish between specular and rough diffuse surfaces. Instead, we propose an environmental map with a single pronounced light source, a weaker ambient lighting component, and additional augmentations with horizontal rotations. For final evaluation, we also employ a distinct studio lighting environment, different from the one used for training (for more details, see Appendix A).
To disentangle the effects of the model from the lighting strategy, we evaluated all combinations of training and evaluation setups against Paint-It in Table 4. Our full pipeline (CasTex with our proposed envlight) consistently achieves the best metrics across both evaluation scenarios, demonstrating its robust performance. Even when both models are trained and evaluated on the same, less expressive Paint-It environment map, CasTex outperforms Paint-It, highlighting an advantage of our model that stems from its design.
Res. Method Runtime VRAM FID KID
(min, $\downarrow$) (GB, $\downarrow$) ($\downarrow$) ($\times 10^{-3}$, $\downarrow$)
1K CasTex (DIP) $\approx 14$ 20.7 20.8 3.7
CasTex $\mathbf{\approx 12}$ 10.5 19.5 2.4
2K CasTex (DIP) $\approx 24$ 52.3 21.1 3.9
CasTex $\mathbf{\approx 12}$ 10.8 19.7 2.6
4K CasTex (DIP) OOM OOM
CasTex $\mathbf{\approx 12}$ 12 19.6 2.6
[…]
DIP versus explicit parameterization. Following the approach of our closest competitor, Paint-it, we integrate Deep Image Prior (DIP) [35] into our method and evaluate its impact on texture optimization (see Section 4 for details). We test our base model CasTex (XL+L) against its modification with implicit texture representation, CasTex (DIP, XL+L). Table 5 shows the results for quality, performance, and scalability across multiple resolutions.
The analysis reveals two key findings. First, the DIP-based variant does not provide a significant quality benefit while being noticeably slower at all resolutions. Second, our method with explicit parameterization demonstrates remarkable scalability, maintaining a nearly constant runtime and only marginal VRAM growth. This confirms that for our cascaded approach, the added complexity of DIP is unnecessary for quality and scalability, making our explicit parameterization the best choice for practical applications.